OpenStudy (anonymous):

Bernard is realizing that his dimensions are not fitting the way that he wants it to because he forgot about the deck. The deck is going to be a rectangular space that borders the pool. Also, for the pool to fit just right there must be a special product represented when either the area or volume expressions are determined. He wants you to design your own rectangular pool using polynomial expressions, including the deck to borders the pool.

OpenStudy (anonymous):

A. Include a diagram or sketch of your pool and deck. Your diagram must have all dimensions represented with polynomials and ensure that a special product will be used. B.Determine the polynomial expression to represent the area of land space that the pool will cover. C.Determine the area of the deck that borders the pool based on your dimensions.

OpenStudy (anonymous):

x represents the width of the pool the length is 3*x+3=3xx+3 and the depth is 2*length -7 = 2(3x+3)-7 = 6x-1